# Stepwise Guided Policy Optimization: Coloring your Incorrect Reasoning in GRPO

### v4

@@ 3.2 Accelerating Learning Dynamics
Suppose that $p_{\texttt{GRPO}}^{(0)}=q_{\texttt{GRPO}}^{(0)}=p_{\texttt{SPO}}^{(0)}=q_{\texttt{SPO}}^{(0)}=\frac{1}{2}$ and $\eta=1$ [^3] We use the unit stepsize for simplicity. Our results are valid for any sufficiently small step size. for GRPO and SGPO. Then, we have that (i) GRPO and SGPO achieve successful learning: $p_{\texttt{GRPO}}^{(k)},q_{\texttt{GRPO}}^{(k)},p_{\texttt{SGPO}}^{(k)},q_{\texttt{SGPO}}^{(k)}\to 1$ as $k\to+\infty$; (ii) SGPO outperforms GRPO in learning the “good” action in the first step: $p_{\texttt{SGPO}}^{(k)}>p_{\texttt{GRPO}}^{(k)}$ for all $k\geq 1$; (iii) SGPO outperforms GRPO in learning the optimal policy: $p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}>p_{\texttt{GRPO}}^{(k)}q_{\texttt{GRPO}}^{(k)}$ for all $k\geq 1$.
Proof Sketch. For (i), we first show the sequence $(p_{\texttt{SGPO}}^{(k)})_{k\geq 1}$ is strictly increasing and bounded in $(0,1)$ (see Lemmas C.4(i) and C.4(ii)). Thus, it converge to some value $c\in(0,1]$. Take limit as $k\to\infty$ on both sides of the update rule, we obtain the only feasible solution $c=1$. Similarly, we show $q_{\texttt{SGPO}}^{(k)},p_{\texttt{GRPO}}^{(k)},q_{\texttt{GRPO}}^{(k)}\to 1$ as $k\to\infty$. For (ii), we use induction. For the base case, we use the update rule and Lemma C.1(ii). Assume $p_{\texttt{SGPO}}^{(k)}>p_{\texttt{GRPO}}^{(k)}$ for some $k\geq 1$, we derive $p_{\texttt{SGPO}}^{(k+1)}>p_{\texttt{GRPO}}^{(k+1)}$ using Lemmas C.1(i) and C.1(ii). For (iii), we show $p_{\texttt{GRPO}}^{(k)}=q_{\texttt{GRPO}}^{(k)}$ for all $k\geq 1$ using induction. It suffices to show $p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}>(p_{\texttt{GRPO}}^{(k)})^{2}$ for all $k\geq 1$. We prove using induction again. For the base case, we use Lemma C.1(iv). It remains to show that $p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}>(p_{\texttt{GRPO}}^{(k)})^{2}$ implies $p_{\texttt{SGPO}}^{(k+1)}q_{\texttt{SGPO}}^{(k+1)}>(p_{\texttt{GRPO}}^{(k+1)})^{2}$. Note that Lemma C.4(iii) guarantees $p_{\texttt{SGPO}}^{(k)}>q_{\texttt{SGPO}}^{(k)}$ for all $k\geq 1$. Thus, Lemma C.2 implies that
$p_{\texttt{SGPO}}^{(k+1)}q_{\texttt{SGPO}}^{(k+1)}=\tfrac{1}{A(p_{\texttt{SGPO}}^{(k)})B(p_{\texttt{SGPO}}^{(k)},q_{\texttt{SGPO}}^{(k)})}>\tfrac{1}{C\left(\sqrt{p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}}\right)^{2}},$
where the functions $A(\cdot)$, $B(\cdot)$ and $C(\cdot)$ are defined in Lemma C.2. By using Lemma C.1(iii), we complete the induction by applying our induction hypothesis. $\Box$
@@
AMC23
avg@16
AIME24
avg@16
MATH500
pass@1
Olympiads
pass@1
LIMO
pass@1
@@
Theorem 3.3 presents one of the first theoretical analyses of GRPO with multiple samples and multi-step reasoning in the context of LLM reasoning. The first part establishes that SGPO converges to the optimal policy. The second and third parts demonstrate that SGPO both accelerates the acquisition of partially correct reasoning steps and preserves partial reasoning ability even when the final answer is incorrect. Importantly, the theorem provides a per-iteration comparison of learning under different reward mechanisms – an aspect rarely examined in previous works. The provable improvement in learning the optimal policy is also consistent with our numerical findings as partly shown in Figure 1; see Section C.4 for further details.
@@ 4 Experiments
We demonstrate the benefits of differentiating negative samples through experiments in both offline and online settings. Offline RL is more computationally efficient, offering faster training, reduced memory consumption, and improved stability. In contrast, online RL provides greater flexibility and learning capacity, and has become the standard approach in large-scale reasoning models such as DeepSeek-R1 (Guo et al., 2025a).
@@ 4.1 Offline Training
For baselines, we consider strong models without further fine-tuned on math-specific SFT datasets, namely Qwen2.5-14B-Instruct and Qwen2.5-32B-Instruct. Prior work showed that a small set of carefully curated prompts significantly enhance the reasoning capability. Accordingly, we adopt the GAIR/LIMO dataset (Ye et al., 2025) as the training set, which has demonstrated strong potential for improving the reasoning performance of large-scale (32B) models in offline SFT. Evaluation is conducted on four standard math reasoning benchmarks: AIME24, AMC23, MATH500 (Hendrycks et al., 2021), and OlympiadBench (He et al., 2024). Our aim is to highlight the rich learning signal contained in all-negative-sample groups, showing that training exclusively on them can still yield performance gains. For benchmarks with fewer than 100 questions (AMC23, AIME24), we report avg@16 results with a decoding temperature of $0.6$ and $\texttt{Top\_P}=0.95$. For benchmarks with more than 100 questions, we report pass@1 results using greedy decoding. Across all experiments, the maximum decoding length is set to $32768$ tokens.
@@
We conduct offline RL training to demonstrate that utilizing all-negative-sample groups can enhance the reasoning abilities of LLMs. For comparison, we also include positive-only offline RL training. As shown in Table 1, SGPO with negative samples consistently improve performance across most of benchmarks, in some cases even surpassing models trained solely on positive samples. Notably, in the 14B model experiment, training on negative samples yields improvements on four benchmarks relative to the positive-sample baseline. These findings underscore the utility of negative samples, which should not be discarded in online GRPO training; see more discussions in Section 4.4.
@@ 4.2 Online Training
For baseline models, we consider applying Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, QwQ-32B, DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B. Online GRPO training is implemented using the verl framework (Sheng et al., 2025). For the step-wise judge model, we adopt a diverse set of LLMs, ranging from closed-source models with strong reasoning capabilities (o4-mini, Claude3.7) to open-source models that are accessible to the community, including DeepSeek-V3-0324, Qwen3-235B-A22B, and QwQ-32B.
@@
For GRPO training, we use the AIME collections from 1997 to 2023 provided in DeepScaler (Luo et al., 2025b), training for 12 epochs. All training questions are in English, while evaluation benchmarks include multilingual questions. Notably, negative samples learned during training generalize well to out-of-domain mathematical reasoning tasks. SGPO training follows the same setup. With batch-simultaneous processing, judge model calls take  90 seconds per batch of negatives, adding  10% wall-clock time relative to rollout and update. Step-wise supervision is applied only to all-negative-sample groups during the first three epochs, as we expect this duration to suffice for the model to internalize corrective signals; beyond this point, unresolved examples are more indicative of model capacity limits than learnability. For all models, rollout length is fixed at $8192$ tokens and group size at $8$. Models less than 8B are trained on 8 H100, 14B models on 16 H100, and 32B models on 32 H200. We adopt the default KL coefficient and learning rate from the verl training script (Sheng et al., 2025), and use the LIMO evaluation script (Ye et al., 2025), both of which are standard practices in the community.
A key insight from Table 2 is that stronger models generate higher-quality negative samples, which substantially aid learning. As model capability improves, so does the informativeness of its mistakes. Negative samples broadly fall into two categories: (i) correct reasoning trajectories truncated by output length limits, and (ii) trajectories containing logical errors. The first type remains highly valuable – yet discarded in GRPO – since it preserves meaningful reasoning steps, motivating our step-wise judge model. The second type, though incorrect, still provides informative signals, especially when all samples fail on genuinely challenging problems. Notably, stronger distilled models average 6K tokens per response, compared to only 1K tokens for weaker base models, making truncated but informative negative samples more common in the stronger case. Likewise, their erroneous responses also tend to be richer and more useful for step-level judgment.
@@
Table 5: Evaluation results are reported in terms of pass@16 across benchmarks. The first two columns show the total number of questions and the number solved within 16 attempts, while the last two columns report the number of unique questions solved by one method but not the other.
@@ 4.3 Other Ablation Studies
To assess reliability of judge models, we evaluate our approach not only with strong closed-source reasoning models but also with publicly available models of weaker capacity: DeepSeek-V3, Qwen3-235B and QwQ-32B. As shown in Table 2 (best-tuned results) and Table 3 (multiple runs with weaker judges), performance remains stable, indicating that weaker judges do not significantly degrade outcomes. We attribute this reliability to two design choices: (i) first-step error identification with a reference answer. SGPO requires the judge only to verify each step against the reference, not to solve the problem, thereby reducing task difficulty and avoiding the pitfalls of generic PRMs; (ii) reward stability parameters $\beta$ and $\gamma$, which filter and smooth noisy signals. As confirmed by ablations, removing $\beta$ and $\gamma$ increases variance and weakens performance. To improve verification, we incorporate a Grok4-Heavy-inspired strategy: multiple independent evaluations by the judge model, with the error position selected by majority voting. Using QwQ-32B as the judge model, DeepSeek-R1-Distill-Qwen-7B as the base model, and four rollouts per judgment, we have observed noticeable gains in consistency and stability (see Table 4).
While avg@16 measures average performance across rollouts, pass@16 reflects the ability to solve new questions with multiple attempts. As shown in Table 5, SGPO’s gains in pass@16 stem directly from leveraging negative samples. Learning only from solvable problems reinforces existing ability, whereas all-negative-sample groups correspond to genuinely difficult questions where GRPO consistently fails. These are precisely the cases where additional feedback can be most valuable. By providing step-level signals, SGPO rewards near-misses by reinforcing correct reasoning up to the first error, penalizes early failures by discouraging persistent error modes, and exposes blind spots by turning hard cases into informative training signals. In this regard, SGPO provides benefits that GRPO cannot match, covering more hard problems and providing sharper credit assignment, which translates to faster and more reliable learning under realistic compute budgets.
@@
By leveraging richer early-stage signals from negative samples, SGPO can achieve faster and stronger performance than GRPO. As shown in Figure 2, SGPO continues improving beyond epoch 5 by solving several additional hard training problems, whereas GRPO plateaus. This improvement stems from informative negative samples that help resolve previously unsolved problems as also shown in Table 5. We also find that SGPO converges more rapidly to a deterministic policy, as illustrated by the training trajectories in Figure 3 (see Appendix B).

### v5

@@ 3.2 Accelerating Learning Dynamics
Suppose that $p_{\texttt{GRPO}}^{(0)}=q_{\texttt{GRPO}}^{(0)}=p_{\texttt{SGPO}}^{(0)}=q_{\texttt{SGPO}}^{(0)}=\frac{1}{2}$ and $\eta=1$ [^3] We use the unit stepsize for simplicity. Our results are valid for any sufficiently small step size. for GRPO and SGPO. Then, we have that (i) GRPO and SGPO achieve successful learning: $p_{\texttt{GRPO}}^{(k)},q_{\texttt{GRPO}}^{(k)},p_{\texttt{SGPO}}^{(k)},q_{\texttt{SGPO}}^{(k)}\to 1$ as $k\to+\infty$; (ii) SGPO outperforms GRPO in learning the “good” action in the first step: $p_{\texttt{SGPO}}^{(k)}>p_{\texttt{GRPO}}^{(k)}$ for all $k\geq 1$; (iii) SGPO outperforms GRPO in learning the optimal policy: $p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}>p_{\texttt{GRPO}}^{(k)}q_{\texttt{GRPO}}^{(k)}$ for all $k\geq 1$.
Proof. To show (i), recall that the sequence $(p_{\texttt{SGPO}}^{(k)})_{k\in\mathbb{N}}$ is strictly increasing and bounded in $(0,1)$ from Lemmas B.4(i) and B.4(ii), so it converges to some value $c\in(0,1]$. Taking limit as $k\to\infty$:
$1=\lim_{k\to\infty}\tfrac{p_{\texttt{SGPO}}^{(k+1)}}{p_{\texttt{SGPO}}^{(k)}}=\lim_{k\to\infty}\tfrac{1}{(1-p_{\texttt{SGPO}}^{(k)})e^{-p_{\texttt{SGPO}}^{(k)}(1-p_{\texttt{SGPO}}^{(k)})}+p_{\texttt{SGPO}}^{(k)}}=\tfrac{1}{(1-c)e^{-c(1-c)}+c}.$
Using the simple Taylor lower bound $e^{-x}\geq 1-x$, we have
To show (ii), consider the base case:
where the inequality follows from Lemma B.1(ii). Thus, we use induction and assume $p_{\texttt{SGPO}}^{(k)}>p_{\texttt{GRPO}}^{(k)}$ for some $k\geq 1$. Then, we have
where the first inequality uses Lemma B.1(i) and the second one uses Lemma B.1(ii). Thus, $p_{\texttt{SGPO}}^{(k+1)}>p_{\texttt{GRPO}}^{(k+1)}$ and the induction is done. We have proved that $p_{\texttt{SGPO}}^{(k)}>p_{\texttt{GRPO}}^{(k)}$ for all $k\geq 1$.
In addition, we recall the update rule of $p_{\texttt{SGPO}}^{(k)}$ and $q_{\texttt{SGPO}}^{(k)}$ as
$\begin{array}[]{lcl}p_{\texttt{SGPO}}^{(k+1)}&=&\tfrac{1}{(1/p_{\texttt{SGPO}}^{(k)}-1)\exp(-p_{\texttt{SGPO}}^{(k)}(1-p_{\texttt{SGPO}}^{(k)}))+1}\\ q_{\texttt{SGPO}}^{(k+1)}&=&\tfrac{1}{(1/q_{\texttt{SGPO}}^{(k)}-1)\exp(-(p_{\texttt{SGPO}}^{(k)})^{2}q_{\texttt{SGPO}}^{(k)}(1-q_{\texttt{SGPO}}^{(k)}))+1}\end{array}.$
It suffices to show $p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}>(p_{\texttt{GRPO}}^{(k)})^{2}$ for all $k\geq 1$. We prove by induction. For the base case,
$\sqrt{p_{\texttt{SGPO}}^{(1)}q_{\texttt{SGPO}}^{(1)}}=\sqrt{\tfrac{1}{1+e^{-1/4}}\cdot\tfrac{1}{1+e^{-1/16}}}>\tfrac{1}{1+e^{-1/8}}=p_{\texttt{GRPO}}^{(1)}.$
The above inequality holds true since Lemma B.1(iv) implies
$2\log(1+e^{-1/8})>\log(1+e^{-1/4})+\log(1+e^{-1/16}),$
It remains to show that $p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}>(p_{\texttt{GRPO}}^{(k)})^{2}$ implies $p_{\texttt{SGPO}}^{(k+1)}q_{\texttt{SGPO}}^{(k+1)}>(p_{\texttt{GRPO}}^{(k+1)})^{2}$ for $k\geq 1$. By Lemma B.4(iii), we know $p_{\texttt{SGPO}}^{(k)}>q_{\texttt{SGPO}}^{(k)}$ for all $k\geq 1$. Thus, Lemma B.2 implies that
Using Lemma B.1(iii), we complete the induction by applying our induction hypothesis:
$\tfrac{1}{C\left(\sqrt{p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}}\right)^{2}}=\left(\exp\left(f_{21}\left(\sqrt{p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}},\sqrt{p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}}\right)\right)\right)^{2}>\left(\exp(f_{21}(p_{\texttt{GRPO}}^{(k)},p_{\texttt{GRPO}}^{(k)}))\right)^{2}=(p_{\texttt{GRPO}}^{(k+1)})^{2}.$
This completes the proof. $\Box$
@@
AMC23 avg@16 AIME24 avg@16 MATH500 pass@1 Olympiads pass@1 LIMO pass@1
@@
Theorem 3.3 presents one of the first theoretical analyses of GRPO with multiple samples and multi-step reasoning in the context of LLM reasoning. The first part establishes that SGPO converges to the optimal policy. The second and third parts demonstrate that SGPO both accelerates the acquisition of partially correct reasoning steps and preserves partial reasoning ability even when the final answer is incorrect. Importantly, the theorem provides a per-iteration comparison of learning under different reward mechanisms – an aspect rarely examined in previous works. The provable improvement in learning the optimal policy is also consistent with our numerical findings. We plot the resulting learning curves of our numerical simulation in Figure 2. The left panel shows the probability of selecting the “good” action in the first step at iteration $k$ (i.e., $p_{\texttt{SGPO}}^{(k)}$ vs. $p_{\texttt{GRPO}}^{(k)}$), while the right panel shows the probability of learning the optimal policy (i.e., $p_{\texttt{SGPO}}^{(k)}q_{\texttt{SGPO}}^{(k)}$ vs. $p_{\texttt{GRPO}}^{(k)}q_{\texttt{GRPO}}^{(k)}$). The results align with the predictions of Theorem 3.3, demonstrating that the likelihood of learning the optimal policy under SGPO consistently exceeds that of GRPO across training.
@@ 4 Experiments
We present the benefits of differentiating negative samples through experiments in both offline and online settings. Offline RL is more computationally efficient, offering faster training, reduced memory consumption, and improved stability. In contrast, online RL provides greater flexibility and learning capacity, and has become the standard approach in large-scale reasoning models such as DeepSeek-R1 (Guo et al., 2025a).
@@ 4.1 Offline Training
For baselines, we consider strong models without further fine-tuned on math-specific SFT datasets, namely Qwen2.5-14B-Instruct and Qwen2.5-32B-Instruct. Prior work showed that a small set of carefully curated prompts significantly enhance the reasoning capability. Accordingly, we adopt the GAIR/LIMO dataset (Ye et al., 2025) as the training set, which has demonstrated strong potential for improving the reasoning performance of large-scale (32B) models in offline SFT. Evaluation is conducted on four standard math reasoning benchmarks: AIME24, AMC23, MATH500 (Hendrycks et al., 2021), and OlympiadBench (He et al., 2024). Our aim is to highlight the rich learning signal contained in all-negative-sample groups, showing that training exclusively on them can still yield performance gains. For benchmarks with fewer than 100 questions (AMC23, AIME24), we report avg@16 results with a decoding temperature of $0.6$ and $\texttt{Top\_P}=0.95$. For benchmarks with more than 100 questions, we report pass@1 results using greedy decoding. Here, pass@k is the percentage of prompts for which at least one of the k sampled responses is correct, while avg@k is the average percentage of corrected samples among the $k$ samples. The maximum decoding length is set to $32768$ tokens.
@@
We conduct offline RL training to demonstrate that utilizing all-negative-sample groups can enhance the reasoning abilities of LLMs. For comparison, we also include positive-only offline RL training. As shown in Table 1, SGPO with negative samples improves average performance and performs competitively on most benchmarks, and in some cases even surpasses models trained solely on positive samples. In particular, in the 14B model experiment, training on negative samples yields improvements on four benchmarks relative to the positive-sample baseline. These findings underscore the utility of negative samples, which should not be discarded in online GRPO training; see further comments in Section 4.4.
@@ 4.2 Online Training
For baselines, we consider applying Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, QwQ-32B, DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B. Online GRPO training is implemented using the verl framework (Sheng et al., 2025). For the step-wise judge model, we adopt a diverse set of LLMs, ranging from closed-source models with strong reasoning capabilities (o4-mini, Claude3.7) to open-source models that are more accessible to the community, including DeepSeek-V3-0324, Qwen3-235B-A22B, and QwQ-32B.
@@
For GRPO training, we use the AIME collections from 1997 to 2023 provided in DeepScaler (Luo et al., 2025b), training for 12 epochs. All training questions are in English, while evaluation benchmarks include multilingual questions. Notably, negative samples learned during training generalize well to out-of-domain mathematical reasoning tasks. SGPO training follows the same setup. With batch-simultaneous processing, judge model calls take  90 seconds per batch of negatives, adding  10% wall-clock time relative to rollout and update. Step-wise supervision is applied only to all-negative-sample groups during the first three epochs, as we expect this duration to suffice for the model to internalize corrective signals; beyond this point, unresolved examples are more indicative of model capacity limits than learnability. Accordingly, the end-to-end wall-clock overhead is upper bounded by approximately $10\%\times 3/12=2.5\%$, and in practice this overhead can be offset by a small reduction of time needed to reach a target performance. For all models, rollout length is fixed at $8192$ tokens and group size at $8$. Models less than 8B are trained on 8 H100, 14B models on 16 H100, and 32B models on 32 H200. We adopt the default KL coefficient and learning rate from the verl training script (Sheng et al., 2025), and use the LIMO evaluation script (Ye et al., 2025), both of which are standard practices in the community. That being said, these benefits are not uniform across every benchmark. When negative samples are short, highly noisy, or dominated by early derailments, step-wise judging provides less actionable signal and can even introduce additional variance through judge noise, so SGPO may match or occasionally underperform GRPO or the baseline on specific tasks. This pattern is consistent with our empirical results and motivates a more nuanced view: SGPO is most effective when the model’s failures retain informative intermediate structure (e.g., truncated near-correct trajectories or localized errors), rather than unstructured breakdowns.
A key insight from Table 2 is that stronger models generate higher-quality negative samples, which aid learning. As model capability improves, so does the informativeness of its mistakes. Negative samples fall into two categories: (i) correct reasoning trajectories truncated by output length limits, and (ii) trajectories containing logical errors. The first type remains highly valuable – yet discarded in GRPO – since it preserves meaningful reasoning steps, motivating our step-wise judge model. The second type, though incorrect, still provides informative signals, especially when all samples fail on genuinely challenging problems. Notably, stronger distilled models average 6K tokens per response, compared to only 1K tokens for weaker base models, making truncated but informative negative samples more common in the stronger case. Likewise, their erroneous responses tend to be richer and more useful for step-level judgment.
@@
Figure 3: Mean and standard deviation over multiple runs across different judge model setups.
@@
Table 5: Evaluation results are reported in terms of pass@16 across benchmarks. The first two columns show the total number of questions and the number solved within 16 attempts, while the last two columns report the number of questions solved by SGPO but not by GRPO ($\text{SGPO}\setminus\text{GRPO}$), and vice versa ($\text{GRPO}\setminus\text{SGPO}$).
@@ 4.3 Other Ablation Studies
To assess reliability of judge models, we evaluate our approach not only with strong closed-source reasoning models but also with publicly available models of weaker capacity: DeepSeek-V3, Qwen3-235B and QwQ-32B. As shown in Table 2 (best-tuned results) and Table 3 (multiple runs with weaker judges), performance remains stable, indicating that weaker judges do not significantly degrade outcomes. We attribute this reliability to two design choices: (i) first-step error identification with the reference answer. SGPO requires the judge only to verify each step against the reference, not to solve the problem, thereby reducing task difficulty and avoiding the pitfalls of generic PRMs; (ii) reward stability parameters $\beta$ and $\gamma$, which set the update inertia and reduce sensitivity to rewards from earlier failed rollouts. As confirmed by ablations, removing $\beta$ and $\gamma$ increases variance and weakens performance. To improve verification, we incorporate a Grok4-Heavy-inspired strategy: multiple independent evaluations by the judge model, with the error position selected by majority voting. Using QwQ-32B as the judge model, DeepSeek-R1-Distill-Qwen-7B as the base model, and four rollouts per judgment, we observed noticeable gains in consistency and stability (see Table 4). Figure 3 visualizes the aggregated results under different judge modes from Tables 3 and 4.
@@
Although avg@16 measures average performance across rollouts, pass@16 reflects the ability to solve new questions with multiple attempts. As shown in Table 5, SGPO’s gains in pass@16 stem directly from leveraging negative samples. Learning only from solvable problems reinforces existing ability, whereas all-negative-sample groups correspond to genuinely difficult questions where GRPO consistently fails. These are precisely the cases where additional feedback can be most valuable. By providing step-level signals, SGPO rewards near-misses by reinforcing correct reasoning up to the first error, penalizes early failures by discouraging persistent error modes, and exposes blind spots by turning hard cases into informative training signals. In this regard, SGPO can provide additional benefits over GRPO, covering more hard problems and providing sharper credit assignment, which translates to more reliable learning under realistic compute budgets.
By leveraging richer early-stage signals from negative samples, SGPO can achieve competitive performance with GRPO, and in some cases better performance or improved coverage on hard problems. As shown in Figure 4, SGPO continues improving beyond epoch 5 by solving several additional hard training problems, whereas GRPO plateaus. This improvement stems from informative negative samples that help resolve previously unsolved problems as also shown in Table 5. In line with our theoretical finding that SGPO converges faster than GRPO, empirical metrics offer supporting evidence. Prior work on RLVR entropy highlights its link to performance: Cui et al. (2025) showed that lower policy entropy under correct signals correlates with stronger policies, while Agarwal et al. (2025) demonstrated that directly minimizing entropy can improve performance. As shown in Figure 5, SGPO reduces policy entropy more rapidly than GRPO, indicating faster convergence toward deterministic RLVR behavior with higher rollout confidence. This matches our theoretical results, confirming that step-wise signals accelerate convergence.